exercise-Assume Mr mulugeta had budget of 150$ for the
Consumption of tomato (X)and Potatol(y) Per month.
assume Currently Price of tomato is 50 $Per kg
& Price of Potato is 10 $ per kg. It Wility function
Of Mr mulugeta is given by Tu=5yx^{2} find the
quantities of tomato and Potato that maximizes
his utility ?
1
Expert's answer
2021-09-22T07:51:18-0400

"Tu=5yx^2"

Budget line "M = P(x) x +P(y)y"

"150 = 50x + 10y"

At optimum :

"\frac{MU(x)}{P(x)}= \frac{MU(y)}{P(y)} \\

MU(x) = \frac{ΔU}{Δx} = 10xy \\

MU(y) = \frac{ΔU}{Δy} = 5x^2 \\

\frac{10xy}{50} = \frac{5x^2}{10} \\

\frac{y}{5} = \frac{x}{2} \\

y=2.5x"

Putting y=2.5x in budget line

"150 = 50 x + 10 \times 2.5x \\

150 = 75x \\

x = \frac{150}{75}=2 \\

y= 5"